OpenStudy (anonymous):

Solve: 8(x − 1) = 6(x + 2) − 4 0 8 - 1/2 3

OpenStudy (anonymous):

8

OpenStudy (anonymous):

8x-8=6x+12-4 8x-6x=12+8-4 2x=16 x=8